DVCS DAQ Oct 18th

From Hall A Wiki
Jump to: navigation, search

Status :

  • Top crate works ok
  • move to other crate improved : can take full calorimeter data
  • trigger read out, firmware being checked - test during access
  • data with issues but most likely good enough for elastic calibration
  • still need to validate some components for cross sections measurement


Issues :

  • ARS empty FIFO message and synchronization [[1]]
  • data still shifts in the buffer
  • ARS initialization - not clear if we start shifted by one
  • spares : no spare left - one crashing - one loosing synch

Requirements for production :

  • now
    • check ARS mapping with calorimeter
    • Dead time calculation
    • sync check
    • helicity
    • ARS trigger efficiency - data taking efficiency
    • calorimeter calibration
  • short term
    • trigger
    • logic module with triggers and bit pattern, on board scaler for dead time

Data to be analyzed :

  • elastic data
  • sync data
  • optics data
  • BCM calibration
  • Pion rejector calibration


Possible plan :

  • today
    • During access :
      • check trigger
      • test sync checks, scalers, dead time
      • trigger power cycling
      • EPICS crate - HV cratea power cycling
      • ARS mapping check with LED
    • Meeting with DAQ/ Electronics group at 2 PM room F227
  • continue elastic calibration Monday owl, update ARS firmware
  • Tuesday, Wednesday - trigger commissioning
  • polarity change Thursday - take DVCS production single and coincidence - check consistency between S2 trigger and S2 trigger in logic module - Question stay at 5 pass or go to lower pass ?